# Relative Gieseker’s problem on $F$-divided bundles

### v1

@@ Abstract
Let $f:X\to Y$ be a surjective morphism of normal projective varieties defined over an algebraically closed field of positive characteristic. We prove that if the induced map on étale fundamental groups is surjective then the corresponding map on $F$-divided fundamental groups is faithfully flat. We also prove an analogous result for isomorphisms. This generalizes and strengthens a recent result of X. Sun and L. Zhang [SZ25], which in turn generalized earlier results of H. Esnault and V. Mehta [EM10] and I. Biswas, M. Kumar, and A. J. Parameswaran [BKP25]. An important new ingredient in our proof is an analogue of B. Bhatt’s and P. Scholze’s descent theorem [BS17, Theorem 1.3] for $F$-divided bundles.
@@ Introduction
Let $X$ be a connected scheme of finite type over some algebraically closed field $k$. In [Gro68, Appendix] A. Grothendieck defined a coherent stratified sheaf on $X/k$ as a coherent sheaf $E$ on $X$ together with an isomorphism $\varphi:p_{1}^{*}E\xrightarrow{\simeq}p_{2}^{*}E$ on the formal completion $\widehat{X\times_{k}X}$ of $X\times_{k}X$ along the diagonal, subject to the cocycle condition $p_{23}^{*}(\varphi)\circ p_{12}^{*}(\varphi)=p_{13}^{*}(\varphi)$ on the formal completion     $X\times_{k}X\times_{k}X$ $\kern-0.6pt\bigwedge\kern-0.6pt$      of $X\times_{k}X\times_{k}X$ along the diagonal. Such stratified sheaves can also be viewed as crystals on the infinitesimal site $(X/k)_{\sf inf}$. In a later work [Gro70], Grothendieck used stratified sheaves to interpret the Riemann-Hilbert correspondence for proper complex varieties. Namely, if $X$ is a proper complex variety then the category of representations of the topological fundamental group $\pi_{1}^{\sf top}(X^{\sf an})$ of the analytification of $X$ is equivalent to the category ${\sf Coh}^{\sf strat}(X)$ of coherent stratified sheaves on $X/{\mathbb{C}}$ (see [Gro70, (4.1.5)]). In particular, ${\sf Coh}^{\sf strat}(X)$ is equivalent to the category ${\sf Vect}^{\sf strat}(X)$ of stratified vector bundles on $X/{\mathbb{C}}$. Later N. Saavedra Rivano in [SR72, Chapitre VI, 1.2] proved that in general the category ${\sf Vect}^{\sf strat}(X)$ is Tannakian and hence when fixing a $k$-point of $X$, Tannakian duality gives us a stratified fundamental group $\pi^{\sf strat}_{1}(X)$. This is a purely algebraic object, and for complex varieties its representations are equivalent to that of $\pi_{1}^{\sf top}(X^{\sf an})$. Using this language we can then reformulate (and slightly strengthen) [Gro70, Theoreme 4.2] as follows:
@@
If $k$ has positive characteristic then the category ${\sf Vect}^{\sf strat}(X)$ of stratified vector bundles on $X/k$ is equivalent to the category ${\sf Vect}^{\sf perf}(X)$ of so called $F$-divided bundles (see [ES16, Proposition 3.4]). An $F$-divided vector bundle on $X$ is a sequence ${\mathbb{E}}:=\{E_{i},\sigma_{i}\}_{i\in{\mathbb{Z}}_{\geq 0}}$ of vector bundles $E_{i}$ on $X$ and ${\mathcal{O}}_{X}$-isomorphisms $\sigma_{i}:F_{X}^{*}E_{i+1}\xrightarrow{\simeq}E_{i}$. Again if $X$ is smooth then giving a stratified vector bundle is equivalent to giving a $D$-module (see Katz’s [Gie75, Theorem 1.3]). However, this condition is significantly stronger than merely specifying a vector bundle with an integrable connection.
As in the characteristic zero case, ${\sf Vect}^{\sf perf}(X)$ is a Tannakian category and when fixing a $k$-point, Tannakian duality gives rise to an $F$-divided fundamental group $\pi^{F\sf{\textendash div}}_{1}(X)$. Then we have the following analogue of Grothendieck’s Theorem 0.1 (see Theorem 3.8 and Theorem 3.11).
@@
Let $f:X\to Y$ be a surjective morphism of normal projective varieties defined over an algebraically closed field of positive characteristic.
@@
    If $Y$ is smooth and $f_{*}:\pi^{\sf{\acute{e}t}}_{1}(X)\to\pi^{\sf{\acute{e}t}}_{1}(Y)$ is an isomorphism then $f_{*}:\pi^{F\sf{\textendash div}}_{1}(X)\to\pi^{F\sf{\textendash div}}_{1}(Y)$ is an isomorphism.
The absolute version of the above theorem when $X$ is smooth and $Y$ is a point was proven in [EM10] in response to a problem posed by Gieseker (see [Gie75, §2]). Recently, this result was generalized to normal proper varieties in [LZ25]. Theorem 0.2 positively answers [Esn13, Question 3.2 (i)] in the case of a morphism between normal projective varieties. Note that for our proof it is essential to consider normal varieties even if one is interested only in smooth varieties as in the original formulation of the question.
The first part of the above theorem was previously unknown. The second part, in the case where both $X$ and $Y$ are smooth, appears in [SZ25], who provide a different and more involved proof. In the special case where $f$ is a finite genuinely ramified morphism, the second part of the above theorem was proven for smooth varieties in [BKP25] and this result plays a crucial role in our proof. This is the only reason why we add the assumption on smoothness of $Y$ in the second part of Theorem 0.2. Unfortunately, in the proofs of both [BKP25, Proposition 4.1] and [SZ25, Lemma 3.2], the smoothness of $Y$ is essential, and the behaviour of étale fundamental groups under spreading out of a finite morphism between general normal varieties is not well understood. However, this assumption is not needed if $f$ has geometrically connected fibers:
@@
The first part of the above theorem follows from more general Corollary 2.5. It is an analogue of [SGA03, Exposé IX, Corollaire 5.6] for the $F$-divided fundamental group. This result was known only in case $f$ is a smooth morphism between smooth projective varieties (see [dS15, Theorem 1.1]). The second part follows from Theorem 3.11 and Lemma 1.3.
One of the main ingredients in our proof of Theorems 0.2 (see Theorem 2.2) and Theorem 0.3 is the following analogue of the descent theorem [BS17, Theorem 1.3] of B. Bhatt and P. Scholze in the case of $F$-divided bundles, which might be of independent interest.
@@
In fact, [BS17, Theorem 1.3] already implies that we can descend each $E_{i}$ separately after some finite purely inseparable map. Since such maps are dominated by some iterates of the absolute Frobenius, there exists some integer $e\geq 0$ such that $(F_{X}^{e})^{*}E_{0}$ descends to $Y$. Then, for all $i\geq 0$, the bundles $(F_{X}^{e+i})^{*}E_{i}$ descend to $Y$, so [BS17, Theorem 1.3] does not give us any additional information. In particular, it does not allow us to descend $\mathbb{E}$} to $Y$. The main challenge in our proof is to find a single map that works for all $E_{i}$ at the same time and notably, our argument does not rely on the result of Bhatt and Scholze.
The proof of the first part of Theorem 0.2 uses just Theorem 0.4 and various results on genuinely ramified morphism. But as in all the previous papers on the subject, the proof of the second part of Theorem 0.2 depends on important Hrushovski’s theorem [Hru04, Corollary 1.2] and boundedness of families of semistable sheaves with fixed numerical invariants (see [Lan04, Theorem 4.4]). In this paper, this last result is hidden in X. Sun’s generalization of Simpson’s representation spaces to positive characteristic (see [Sun19, Theorem 2.3]), which is also an important ingredient of the proof. As in [ES19] one could also try to prove Theorem 0.2, (2) using moduli spaces of semistable vector bundles constructed in [Lan04]. Then it is easy to descend simple objects of ${\sf Vect}^{\sf perf}(X)$ but as in [ES19, Section 5] this approach runs into technical difficulties involving study of universal extensions on products of moduli spaces making it impractical due to a horrible notation and complications.
@@ 1.1 $F$-divided bundles
Below we recall the following result from [LZ25] that generalizes J.-P.-S. dos Santos’s result from [dS07] (see also [ES16, Proposition 3.3]).
@@
We will also need the following lemma from [LZ25, Lemma 5.9].
@@
In addition, we require the following simple result (see [ES16, Proposition 3.3] and [LZ25, Lemma 2.3]).
@@ 1.2 Representation schemes
In the proof of Theorem 3.11 we use existence of a certain moduli scheme of vector bundles that we recall in this subsection. Let us fix a positive integer $r$, a projective morphism $f:X_{S}\to S$ of schemes of finite type over some fixed field and an $f$-very ample line bundle ${\mathcal{O}}_{X}(1)$. Let us assume that the fibers of $f$ are geometrically connected and there exists a section $x_{S}:S\to X_{S}$ of the morphism $f$. The following theorem was proven by X. Sun in [Sun19, Theorem 2.3]. We formulate only a special case that is more suitable for our applications.
@@
The above version is slightly different to the one constructed in [Sun19] and then used in [BKP25] and [SZ25]. It corresponds to an open subscheme of the corresponding representation scheme for the Hilbert polynomial of the trivial bundle, classifying vector bundles (this is an open condition, e.g., by [HL10, Lemma 2.1.8]) with numerically vanishing Chern classes (which is also an open condition). This scheme is much better behaved when considering a morphism $f:X\to Y$ and the rational map on representation schemes defined by the pullback. In general, the author does not know a good reason for the behaviour of Hilbert polynomials as claimed in all the above cited papers, e.g., [Sun19, proof of Theorem 4.2], [BKP25, Proposition 2.1] and various places in [SZ25]. In fact, this seems to cause problems even when dealing with the Verschiebung map defined by the Frobenius morphism and which appears all over these papers.
@@ 2 $F$-divided bundles on fibrations
The following theorem is an analogue of [BS17, Theorem 1.3].
@@
By [LZ25, Lemma 1.6] it is sufficient to show that for any ${\mathbb{E}}\in{\sf Vect}^{\sf perf}(Y)$, any subobject ${\mathbb{F}}\subseteq f^{*}{\mathbb{E}}$ is contained in the essential image of $f^{*}$. Thanks to Lemma 2.1, it is enough to show that ${\mathbb{F}}|_{X_{\bar{y}}}$ is trivial for all geometric points $\bar{y}$ of $Y$. Since ${\sf Vect}^{\sf perf}(X_{\bar{y}})$ is a $\kappa(\bar{y})$-Tannakian category (cf. Theorem 1.1) and ${\mathbb{F}}|_{X_{\bar{y}}}$ is a subobject of the trivial object ${\mathbb{E}}|_{X_{\bar{y}}}$, it must be trivial. ∎
@@ 3.1 Genuinely ramified morphisms
The following definition was first introduced in [PS10, Definition 3.1] in the curve case and then generalized to higher dimensional smooth varieties in [BKP25]. In the definition we allow normal varieties.
@@
Let $f:X\to Y$ be a finite morphism of normal projective $k$-varieties. We say that $f$ is genuinely ramified if it is surjective, the field extension $K(X)/K(Y)$ is separable and the induced map $f_{*}:\pi^{\sf{\acute{e}t}}_{1}(X)\to\pi^{\sf{\acute{e}t}}_{1}(Y)$ is surjective.
The proof of the following lemma is the same as that of [BKP25, Lemma 6.4] and we skip it.
@@
Let $f:X\to Y$ be a genuinely ramified finite morphism of normal projective $k$-varieties. Assume that the field extension $K(X)/K(Y)$ is Galois of degree $d$. Then there exists a big open subset $V\subset Y_{\sf reg}$ with $U:=f^{-1}(V)\subset X_{\sf reg}$ and line bundles $L_{j}\subsetneq{\mathcal{O}}_{U}$ for j=1,,d1j=1,...,d-1 such that
@@
As a corollary (see [BKP25, Lemma 6.5] for a slightly weaker version for Gieseker semistable vector bundles in higher dimensions) we obtain the following result:
@@
Let $f:X\to Y$ be a genuinely ramified finite morphism of normal projective $k$-varieties. Then for any slope semistable reflexive sheaf $E$ on $Y$ we have
@@
In the following if $E$ is a reflexive sheaf on $Y$ then we write $f^{[*]}E$ to denote the reflexivization of $f^{*}E$. The above corollary implies the following corollary (see [PS10, Lemma 3.5] for the curve case and [BKP25, Proposition 2.3] for vector bundles on smooth higher dimensional varieties).
@@
Let $f:X\to Y$ be a genuinely ramified finite morphism of normal projective $k$-varieties. Then for any slope semistable reflexive sheaves $E_{1},E_{2}$ on $Y$ with the same slope the natural map
@@
There exists a big open subset $V\subset Y_{\sf reg}$ such that $U:=f^{-1}(V)\subset X_{\sf reg}$ and restrictions of $E_{1}$ and $E_{2}$ to $V$ are locally free. Then the same arguments as that in [BKP25, Proposition 2.3] show that
@@
We will also need the following analogue of [PS10, Lemma 3.5, (b)]:
@@
Let $f:X\to Y$ be a genuinely ramified finite morphism of normal projective $k$-varieties. Then for any slope stable reflexive sheaf $E$ on $Y$ the reflexivized pullback $f^{[*]}E$ is also slope stable.
Proof.
Let $L$ be a splitting field of the function field $K(X)$ over $K(Y)$ and let $Z$ be the normalization of $X$ in $L$. Then we have finite morphisms $Z\to X\to Y$ and it sufficient to prove that the reflexivized pullback of $E$ to $Z$ is stable. So in the following we can assume that the extension $K(X)/K(Y)$ is Galois with a finite Galois group $G$. Now let us consider the socle $S$ of $f^{[*]}E$. Because of its uniqueness $S$ is invariant unfer the action of $G$ and hence it descends to a subsheaf of $E$. Since $E$ is stable this implies that $S=f^{[*]}E$. Since $S$ is a direct sum of slope stable sheaves and we have $k={\mathop{{\rm End\,}}}E\simeq{\mathop{{\rm End\,}}}f^{[*]}E$ by Corollary 3.4, this implies that $f^{[*]}E$ is slope stable. ∎
Using the above facts the following theorem can be proven by adapting the arguments from the proof of [BKP25, Theorem 1.1] dealing with smooth varieties. We omit the proof. It should be noted, however, that the proof of [BKP25, Theorem 1.1] contains an error in the claim that one can spread out infinitely many bundles simultaneously. This issue can be easily corrected by constructing a model for the Zariski closure of the framed bundles in the representation scheme (see an application of this technique below in the proof of Theorem 3.8).
Theorem 3.6.
Let $f:X\to Y$ be a genuinely ramified finite morphism of normal projective $k$-varieties. If $Y$ is smooth and $f_{*}:\pi^{\sf{\acute{e}t}}_{1}(X)\to\pi^{\sf{\acute{e}t}}_{1}(Y)$ is an isomorphism then $f_{*}:\pi^{F\sf{\textendash div}}_{1}(X)\to\pi^{F\sf{\textendash div}}_{1}(Y)$ is an isomorphism.
Unfortunately, [BKP25, Proposition 4.1] that is used in the above theorem uses the Zariski–Nagata purity theorem and hence it can be generalized only in case $Y$ is smooth (see [SZ25, Lemma 3.2]). The use of [SZ25, Lemma 3.2] is the only reason why we add assumption that $Y$ is smooth in the above theorem.

### v2

@@ Abstract
Let $f:X\to Y$ be a proper surjective morphism of varieties defined over an algebraically closed field of positive characteristic. We prove that if $f$ has geometrically connected fibers then the induced homomorphism of $F$-divided fundamental groups is faithfully flat. An important new ingredient in our proof is an analogue of B. Bhatt’s and P. Scholze’s descent theorem [2, Theorem 1.3] for $F$-divided bundles.
As a corollary, we prove that in general if $X$ is normal, $Y$ is smooth, both $X$ and $Y$ are projective, and the induced map on étale fundamental groups is surjective, then the corresponding homomorphism on $F$-divided fundamental groups is faithfully flat. We also establish an analogous result for isomorphisms. This generalizes and strengthens a recent result of X. Sun and L. Zhang [28], which in turn generalized earlier results of H. Esnault and V. Mehta [9] and I. Biswas, M. Kumar, and A. J. Parameswaran [3].
@@ Introduction
Let $X$ be a connected scheme of finite type over some algebraically closed field $k$. In [16, Appendix] A. Grothendieck defined a coherent stratified sheaf on $X/k$ as a coherent sheaf $E$ on $X$ together with an isomorphism $\varphi:p_{1}^{*}E\xrightarrow{\simeq}p_{2}^{*}E$ on the formal completion $\widehat{X\times_{k}X}$ of $X\times_{k}X$ along the diagonal, subject to the cocycle condition $p_{23}^{*}(\varphi)\circ p_{12}^{*}(\varphi)=p_{13}^{*}(\varphi)$ on the formal completion     $X\times_{k}X\times_{k}X$ $\kern-0.6pt\bigwedge\kern-0.6pt$      of $X\times_{k}X\times_{k}X$ along the diagonal. Such stratified sheaves can also be viewed as crystals on the infinitesimal site $(X/k)_{\sf inf}$. In a later work [17], Grothendieck used stratified sheaves to interpret the Riemann-Hilbert correspondence for proper complex varieties. Namely, if $X$ is a proper complex variety then the category of representations of the topological fundamental group $\pi_{1}^{\sf top}(X^{\sf an})$ of the analytification of $X$ is equivalent to the category ${\sf Coh}^{\sf strat}(X)$ of coherent stratified sheaves on $X/{\mathbb{C}}$ (see [17, (4.1.5)]). In particular, ${\sf Coh}^{\sf strat}(X)$ is equivalent to the category ${\sf Vect}^{\sf strat}(X)$ of stratified vector bundles on $X/{\mathbb{C}}$. Later N. Saavedra Rivano in [27, Chapitre VI, 1.2] proved that in general the category ${\sf Vect}^{\sf strat}(X)$ is Tannakian and hence when fixing a $k$-point of $X$, Tannakian duality gives us a stratified fundamental group $\pi^{\sf strat}_{1}(X)$. This is a purely algebraic object, and for complex varieties its representations are equivalent to that of $\pi_{1}^{\sf top}(X^{\sf an})$. Using this language we can then reformulate (and slightly strengthen) [17, Theoreme 4.2] as follows:
@@
If $k$ has positive characteristic then the category ${\sf Vect}^{\sf strat}(X)$ of stratified vector bundles on $X/k$ is equivalent to the category ${\sf Vect}^{\sf perf}(X)$ of so called $F$-divided bundles (see [10, Proposition 3.4]). An $F$-divided vector bundle on $X$ is a sequence ${\mathbb{E}}:=\{E_{i},\sigma_{i}\}_{i\in{\mathbb{Z}}_{\geq 0}}$ of vector bundles $E_{i}$ on $X$ and ${\mathcal{O}}_{X}$-isomorphisms $\sigma_{i}:F_{X}^{*}E_{i+1}\xrightarrow{\simeq}E_{i}$. Again if $X$ is smooth then giving a stratified vector bundle is equivalent to giving a $D$-module (see Katz’s [13, Theorem 1.3]). However, this condition is significantly stronger than merely specifying a vector bundle with an integrable connection.
As in the characteristic zero case, ${\sf Vect}^{\sf perf}(X)$ is a Tannakian category and when fixing a $k$-point, Tannakian duality gives rise to an $F$-divided fundamental group $\pi^{F\sf{\textendash div}}_{1}(X)$. Then we have the following analogue of Grothendieck’s Theorem 0.1 (see Theorem 3.8 and Corollary 3.12).
@@
Let $f:X\to Y$ be a surjective morphism of normal projective varieties defined over an algebraically closed field of positive characteristic. Assume that $Y$ is smooth.
@@
    If $f_{*}:\pi^{\sf{\acute{e}t}}_{1}(X)\to\pi^{\sf{\acute{e}t}}_{1}(Y)$ is an isomorphism then $f_{*}:\pi^{F\sf{\textendash div}}_{1}(X)\to\pi^{F\sf{\textendash div}}_{1}(Y)$ is an isomorphism.
The absolute version of the above theorem when $X$ is smooth and $Y$ is a point was proven in [9] in response to a problem posed by Gieseker (see [13, §2]). Recently, this result was generalized to normal proper varieties in [22]. Theorem 0.2 positively answers [12, Question 3.2 (i)] in the case of a morphism between from normal to smooth projective variety. Note that for our proof it is essential to consider normal varieties even if one is interested only in smooth varieties as in the original formulation of the question.
The first part of the above theorem was previously unknown. The second part, in the case where both $X$ and $Y$ are smooth, appears in [28], who provide a different and more involved proof. In the special case where $f$ is a finite genuinely ramified morphism, the second part of the above theorem was proven for smooth varieties in [3] and this result plays a crucial role in our proof. This is the only reason why we add the assumption on smoothness of $Y$. Unfortunately, in the proofs of both [3, Proposition 4.1] and [28, Lemma 3.2], the smoothness of $Y$ is essential, and the behaviour of étale fundamental groups under spreading out of a finite morphism between general normal varieties is not well understood. However, this assumption is not needed if $f$ has geometrically connected fibers:
@@
The first part of the above theorem follows from more general Corollary 2.5. It is an analogue of [26, Exposé IX, Corollaire 5.6] for the $F$-divided fundamental group. This result was known only in case $f$ is a smooth morphism between smooth projective varieties (see [8, Theorem 1.1]). The second part follows from Theorem 3.11 and Lemma 1.3.
One of the main ingredients in our proof of Theorem 0.2 (see Theorem 2.2) and Theorem 0.3 is the following analogue of the descent theorem [2, Theorem 1.3] of B. Bhatt and P. Scholze in the case of $F$-divided bundles, which might be of independent interest.
@@
In fact, [2, Theorem 1.3] already implies that we can descend each $E_{i}$ separately after some finite purely inseparable map. Since such maps are dominated by some iterates of the absolute Frobenius, there exists some integer $e\geq 0$ such that $(F_{X}^{e})^{*}E_{0}$ descends to $Y$. Then, for all $i\geq 0$, the bundles $(F_{X}^{e+i})^{*}E_{i}$ descend to $Y$, so [2, Theorem 1.3] does not give us any additional information. In particular, it does not allow us to descend $\mathbb{E}$} to $Y$. The main challenge in our proof is to find a single map that works for all $E_{i}$ at the same time and notably, our argument does not rely on the result of Bhatt and Scholze.
The proof of the first part of Theorem 0.2 uses just Theorem 0.4 and various results on genuinely ramified morphism. But as in all the previous papers on the subject, the proof of the second part of Theorem 0.2 depends on important Hrushovski’s theorem [18, Corollary 1.2] and boundedness of families of semistable sheaves with fixed numerical invariants (see [23, Theorem 4.4]). In this paper, this last result is hidden in X. Sun’s generalization of Simpson’s representation spaces to positive characteristic (see [29, Theorem 2.3]), which is also an important ingredient of the proof. As in [11] one could also try to prove Theorem 0.2, (2) using moduli spaces of semistable vector bundles constructed in [23]. Then it is easy to descend simple objects of ${\sf Vect}^{\sf perf}(X)$ but as in [11, Section 5] this approach runs into technical difficulties involving study of universal extensions on products of moduli spaces making it impractical due to a horrible notation and complications.
@@ 1.1 $F$-divided bundles
Below we recall the following result from [22] that generalizes J.-P.-S. dos Santos’s result from [7] (see also [10, Proposition 3.3]).
@@
We will also need the following lemma from [22, Lemma 5.9].
@@
In addition, we require the following simple result (see [10, Proposition 3.3] and [22, Lemma 2.3]).
@@ 1.2 Representation schemes
In the proof of Theorem 3.11 we use existence of a certain moduli scheme of vector bundles that we recall in this subsection. Let us fix a positive integer $r$, a projective morphism $f:X_{S}\to S$ of schemes of finite type over some fixed field and an $f$-very ample line bundle ${\mathcal{O}}_{X}(1)$. Let us assume that the fibers of $f$ are geometrically connected and there exists a section $x_{S}:S\to X_{S}$ of the morphism $f$. The following theorem was proven by X. Sun in [29, Theorem 2.3]. We formulate only a special case that is more suitable for our applications.
@@
The above version is slightly different to the one constructed in [29] and then used in [3] and [28]. It corresponds to an open subscheme of the corresponding representation scheme for the Hilbert polynomial of the trivial bundle, classifying vector bundles (this is an open condition, e.g., by [19, Lemma 2.1.8]) with numerically vanishing Chern classes (which is also an open condition). This scheme is much better behaved when considering a morphism $f:X\to Y$ and the rational map on representation schemes defined by the pullback. In general, the author does not know a good reason for the behaviour of Hilbert polynomials as claimed in all the above cited papers, e.g., [29, proof of Theorem 4.2], [3, Proposition 2.1] and various places in [28]. In fact, this seems to cause problems even when dealing with the Verschiebung map defined by the Frobenius morphism and which appears all over these papers.
1.3 Specialization homomorphism
It is well-known that for a flat proper morphism with geometrically connected fibres the specialization homomorphism of étale fundamental groups need not be surjective in general (see [26, Exposé X, Remarques 2.5]). However, the following lemma shows that this can happen only over a proper closed subset of the base.
Lemma 1.7.
Let $S$ be a locally Noetherian integral scheme and let $f:X\to S$ be a flat proper morphism with geometrically connected fibres. Let $\eta$ be the generic point of $S$ and let $\bar{\eta}$ be a geometric point lying over $\eta$. Then there exists a non-empty open subset $U\subset S$ such that for every point $s\in U$ and a geometric point $\bar{s}$ lying over $s$, the specialization homomorphism ${\sf{sp}}:\pi^{\sf{\acute{e}t}}_{1}(X_{\bar{\eta}})\to\pi^{\sf{\acute{e}t}}_{1}(X_{\bar{s}})$ is surjective.
Proof.
There exists a finite radicial extension $\kappa(\eta)\subset L\subset\kappa(\bar{\eta})$ such that $(X_{L})_{\sf{red}}$ is geometrically reduced (see [14, Proposition 4.6.6]). Let $S^{\prime}\to S$ be a normalization of $S$ in the extension $\kappa(\eta)\subset L$. Let $X^{\prime}$ be the reduced scheme structure on $X\times_{S}S^{\prime}$ and let $f^{\prime}:X^{\prime}\to S^{\prime}$ be the induced map. By [1, Lemma 054Z] the generic fiber of $f^{\prime}$ is reduced and hence it is isomorphic to $(X_{L})_{\sf{red}}$. So by [15, Théorème 12.2.4] there exists a non-empty open subset $U^{\prime}\subset S^{\prime}$ such that the fibers of $f^{\prime}|_{(f^{\prime})^{-1}(U^{\prime})}$ are geometrically reduced. By generic flatness (see [1, Proposition 052B]) we can also assume that $f^{\prime}|_{(f^{\prime})^{-1}(U^{\prime})}$ is flat. Then by [26, Exposé X, Corollaire 2.4] for every geometric point $\bar{s}^{\prime}$ lying over $s^{\prime}\in U^{\prime}$ the specialization homomorphism ${\sf{sp}}:\pi^{\sf{\acute{e}t}}_{1}(X^{\prime}_{\bar{\eta}})\to\pi^{\sf{\acute{e}t}}_{1}(X^{\prime}_{{\bar{s}^{\prime}}})$ is surjective. But if $\bar{s}$ denotes the image of $\bar{s}^{\prime}$ then by [1, Lemma 0C0K] we have a commutative diagram
${\pi^{\sf{\acute{e}t}}_{1}(X^{\prime}_{\bar{\eta}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\sf sp}$${\pi^{\sf{\acute{e}t}}_{1}(X_{\bar{\eta}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\sf sp}$${\pi^{\sf{\acute{e}t}}_{1}(X^{\prime}_{{\bar{s}^{\prime}}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi^{\sf{\acute{e}t}}_{1}(X_{\bar{s}})}$
Since the horizontal maps in the above diagram are isomorphisms, it is sufficient to take as $U$ and non-empty open subset contained in the image of $U^{\prime}\to S$. ∎
@@ 2 $F$-divided bundles on fibrations
The following theorem is an analogue of [2, Theorem 1.3].
@@
By [22, Lemma 1.6] it is sufficient to show that for any ${\mathbb{E}}\in{\sf Vect}^{\sf perf}(Y)$, any subobject ${\mathbb{F}}\subseteq f^{*}{\mathbb{E}}$ is contained in the essential image of $f^{*}$. Thanks to Lemma 2.1, it is enough to show that ${\mathbb{F}}|_{X_{\bar{y}}}$ is trivial for all geometric points $\bar{y}$ of $Y$. Since ${\sf Vect}^{\sf perf}(X_{\bar{y}})$ is a $\kappa(\bar{y})$-Tannakian category (cf. Theorem 1.1) and ${\mathbb{F}}|_{X_{\bar{y}}}$ is a subobject of the trivial object ${\mathbb{E}}|_{X_{\bar{y}}}$, it must be trivial. ∎
@@ 3.1 Genuinely ramified morphisms
The following definition was first introduced in [24, Definition 3.1] in the curve case and then generalized to higher dimensional smooth varieties in [3]. In the definition we allow normal varieties but we restrict to regular parts.
Let $f:X\to Y$ be a finite surjective morphism of normal projective $k$-varieties. Then there exists a big open subset $V\subset Y_{\sf reg}$ with $U:=f^{-1}(V)\subset X_{\sf reg}$. Since we have canonical isomorphisms $\pi^{\sf{\acute{e}t}}_{1}(U)\simeq\pi^{\sf{\acute{e}t}}_{1}(X_{\sf reg})$ and $\pi^{\sf{\acute{e}t}}_{1}(V)\simeq\pi^{\sf{\acute{e}t}}_{1}(Y_{\sf reg})$, the restriction $f|_{U}:U\to V$ defines a canonical homomorphism $f_{*}:\pi^{\sf{\acute{e}t}}_{1}(X_{\sf reg})\to\pi^{\sf{\acute{e}t}}_{1}(Y_{\sf reg})$.
@@
We say that $f$ is regularly genuinely ramified if the field extension $K(X)/K(Y)$ is separable (or, equivalently, $f$ is generically smooth) and the induced homomorphism $f_{*}:\pi^{\sf{\acute{e}t}}_{1}(X_{\sf reg})\to\pi^{\sf{\acute{e}t}}_{1}(Y_{\sf reg})$ is surjective.
The proof of the following lemma is the same as that of [3, Lemma 6.4] and we skip it.
@@
Let $f:X\to Y$ be a regularly genuinely ramified morphism of normal projective $k$-varieties. Assume that the field extension $K(X)/K(Y)$ is Galois of degree $d$. Then there exists a big open subset $V\subset Y_{\sf reg}$ with $U:=f^{-1}(V)\subset X_{\sf reg}$ and line bundles $L_{j}\subsetneq{\mathcal{O}}_{U}$ for j=1,,d1j=1,...,d-1 such that
@@
As a corollary (see [3, Lemma 6.5] for a slightly weaker version for Gieseker semistable vector bundles in higher dimensions) we obtain the following result:
@@
Let $f:X\to Y$ be a regularly genuinely ramified morphism of normal projective $k$-varieties. Then for any slope semistable reflexive sheaf $E$ on $Y$ we have
@@
In the following if $E$ is a reflexive sheaf on $Y$ then we write $f^{[*]}E$ to denote the reflexivization of $f^{*}E$. The above corollary implies the following corollary (see [4, Lemma 4.3] for the curve case and [3, Proposition 2.3] for vector bundles on smooth higher dimensional varieties).
@@
Let $f:X\to Y$ be a regularly genuinely ramified morphism of normal projective $k$-varieties. Then for any slope semistable reflexive sheaves $E_{1},E_{2}$ on $Y$ with the same slope the natural map
@@
There exists a big open subset $V\subset Y_{\sf reg}$ such that $U:=f^{-1}(V)\subset X_{\sf reg}$ and restrictions of $E_{1}$ and $E_{2}$ to $V$ are locally free. Then the same arguments as that in [3, Proposition 2.3] show that
@@
Using the above facts, the following theorem can be proven by adapting the arguments from the proof of [3, Theorem 1.1], which treats the case of smooth varieties. We omit the proof. It should be noted, however, that the proof of [3, Theorem 1.1] contains an error in the claim that one can spread out infinitely many bundles simultaneously. This issue can be easily corrected by constructing a model for the Zariski closure of the framed bundles in the representation scheme (see an application of this technique below in the proof of Theorem 3.11).
Theorem 3.5.
Let $f:X\to Y$ be a finite, generically smooth morphism of normal projective $k$-varieties. If $Y$ is smooth and $f_{*}:\pi^{\sf{\acute{e}t}}_{1}(X)\to\pi^{\sf{\acute{e}t}}_{1}(Y)$ is an isomorphism then $f_{*}:\pi^{F\sf{\textendash div}}_{1}(X)\to\pi^{F\sf{\textendash div}}_{1}(Y)$ is an isomorphism.
Unfortunately, [3, Proposition 4.1] that is used in the above theorem uses the Zariski–Nagata purity theorem and hence it can be generalized only in case $Y$ is smooth (see [28, Lemma 3.2]). The use of [28, Lemma 3.2] is the main reason why we add assumption that $Y$ is smooth in the above theorem. Another reason for assuming that $Y$ is smooth comes from the above lemmas. Note also that one cannot easily prove that in the above setup if $Y$ is smooth and $f_{*}:\pi^{\sf{\acute{e}t}}_{1}(X_{\sf reg})\to\pi^{\sf{\acute{e}t}}_{1}(Y)$ is an isomorphism then $f_{*}:\pi^{F\sf{\textendash div}}_{1}(X_{\sf reg})\to\pi^{F\sf{\textendash div}}_{1}(Y)$ is an isomorphism. In this case one runs into problems with a specialization homomorphism on the étale fundamental group for the non-proper variety $X_{\sf reg}$ (cf. [10, Section 4]).
In the next section we will also need the following analogue of [4, Theorem 4.4]. Since the proof, as in case of Corollary 3.4 is very similar to that from the curve case, we omit it (see also [25, Theorem 1.1]).
@@
Let $f:X\to Y$ be a regularly genuinely ramified finite morphism of normal projective $k$-varieties. Then for any slope stable reflexive sheaf $E$ on $Y$ the reflexivized pullback $f^{[*]}E$ is also slope stable.